Question 2: There are two data classes + and -. The training set is shown in Table 4.2 as follows. Please compute the ent
Answer #1

The standered formula for calculation of entropy is used. The formulas are mentioned in the images below. I have explained all the steps in detail. Furthur clarification can be asked via comments.

Total ho. of positive outputs p = 4 Totd no. of nogetue culpets, h=6 [culpat is the Target Class value] For entropy of ate, h

Here the formula E(A=v) just has two terms as A just takes two distinct values. Like here A=a1 takes two values, that is , T and F. If A takes more values in some other question, then number of terms will be equal to number of distinct vaues taken by A

E[a] = F) = -log(1) - 4 log (45) 00464+00257 - 00721 ECal) = { Pithu. Ela2 = 0) -1 pth VE no. of olistinct values of al. Ea1)

E(a1) is the required weighted average, Again, here also the number of terms in the formula is will be equal to number of distinct vaues taken by A. Here just two ( Same reason as above).

The final answer is 0.767.